What happens to an amount of water when its temperature is reduced from 8°C to 4°C?

Its density increses

its density decreases
its density remains a constant
it vaporizes
Answer #1

Using the equation of density.

\rho =\frac{m}{v}

And using the ideal gas equation

P*v=n*R*T

v=\frac{n*R*T}{P}

If the temperature decreases , the volume decreases . and introducing this value in the equation of density , density tends to increase. Therefore the density increases .